OpenStudy (anonymous):

Multiply and simplify: (2-√7) (2+√7) Answers: -45 4-√7 -3 4-4√7

OpenStudy (anonymous):

@E.ali :))

OpenStudy (psymon):

\[(a-b)(a+b) = a^{2} - b^{2} \]Its kind of a quick way to multiply something specific like this out. So basically a is 2 and b is sqrt(7). a^2 = 2^2 = 4. And b^2 = \[\sqrt{7}*\sqrt{7}=\sqrt{49} = 7\]So this means that a^2 - b^2 = 4 - 7 = -3

OpenStudy (anonymous):

Its C .

OpenStudy (anonymous):

-3